# A diverse, overlooked population of Type Ia supernovae exhibiting mid-infrared signatures of delayed circumstellar interaction

### v1

@@ Appendix B Additional follow-up observations
We acquired Gemini GMOS-N (Hook et al., 2004) spectra on UT 2024-08-06 (SN 2020mvp) and UT 2024-08-07 (SN 2022bbt) using the B480 grating and binning on the Hamamatsu detector, resulting in 1.3 Å spectral resolution, sufficient for distinguishing narrow host lines from the broad or intermediate lines indicative of SN ejecta. The sources were acquired using blind-offsets from nearby stars and each spectrum was integrated for 80 min. The spectra were reduced using standard routines in PypeIt (Prochaska et al., 2020), including flat-fielding and wavelength calibration using arcs. Flux calibration was performed using a standard star observation followed by one-dimensional spectral extraction at the expected position of each SN. No broad or intermediate emission lines were identified in either spectrum (see Section 3.3). For each source, we derived an upper limit on the intermediate emission line flux by rebinning the spectrum to the expected Å width (FWHM km s-1) of the intermediate emission line (Graham et al., 2017), then taking the standard deviation of a Å featureless section of the spectrum near the expected H line.
@@
We observed SN 2020mvp and SN 2021jun in the NIR -band with the WINTER camera (Lourie et al., 2020; Frostig et al., 2024) mounted on the Palomar 1-m telescope on UT 2024-05-12, with a total exposure time of 3600 s for each object. The images were processed using the WINTER data reduction pipeline mirar (Stein et al., 2024) and image subtraction was performed relative to -band images from the UKIRT Hemisphere Survey (Dye et al., 2018). No sources were detected in difference imaging at the location of either SN, with 5 limiting magnitudes of and (AB) respectively.
We also searched for radio emission using the VLA (Condon et al., 1998) on UT 2024-08-28 in C-band, with the array in B-configuration ( beamwidth, to distinguish between host galaxy nuclear emission). For SN 2020mvp, we used 3C286 as a flux calibrator and J1436+2321 as a phase calibrator. For SN 2022bbt, we also used 3C286 as a flux calibrator, while using J1516+1932 as a phase calibrator. The data was calibrated with the VLA CASA Calibration Pipeline using CASA 6.5.4. and imaged manually with the TCLEAN task in CASA, using natural weighting. No radio source was detected for either SN, with a 5 flux density limit of Jy.

### v4

@@ Appendix B Additional follow-up observations
We acquired Gemini GMOS-N (Hook et al., 2004) spectra on UT 2024-08-07 (SN 2022bbt) and Gemini GMOS-S spectra on UT 2024-10-26 (SN 2020yex) using the B480 gratings and binning on the Hamamatsu detectors, resulting in 1.3 Å spectral resolution, sufficient for distinguishing narrow host lines from the broad or intermediate lines indicative of SN ejecta. The sources were acquired using blind-offsets from nearby stars and each spectrum was integrated for 80 min. The spectra were reduced using standard routines in PypeIt (Prochaska et al., 2020), including flat-fielding and wavelength calibration using arcs. Flux calibration was performed using a standard star observation followed by one-dimensional spectral extraction at the expected position of each SN. No broad or intermediate emission lines were identified in either spectrum (see Section 3.3). For each source, we derived an upper limit on the intermediate emission line flux by rebinning the spectrum to the expected Å width (FWHM km s-1) of the intermediate emission line (Graham et al., 2017), then taking the standard deviation of a Å featureless section of the spectrum near the expected H line.
@@
We observed SN 2021jun in the NIR -band with the WINTER camera (Lourie et al., 2020; Frostig et al., 2024) mounted on the Palomar 1-m telescope on UT 2024-05-12, with a total exposure time of 3600 s. The images were processed using the WINTER data reduction pipeline mirar (Stein et al., 2024) and image subtraction was performed relative to -band images from the UKIRT Hemisphere Survey (Dye et al., 2018). No sources were detected in difference imaging at the location of SN 2021jun, with a 5 limiting magnitude of (AB).
We also searched for radio emission from SN 2022bbt using the VLA (Condon et al., 1998) on UT 2024-08-28 in C-band, with the array in B-configuration ( beamwidth, to distinguish between host galaxy nuclear emission). We used 3C286 as a flux calibrator, while using J1516+1932 as a phase calibrator. The data was calibrated with the VLA CASA Calibration Pipeline using CASA 6.5.4. and imaged manually with the TCLEAN task in CASA, using natural weighting. No radio source was detected, with a 5 flux density limit of Jy.